OpenStudy (anonymous):

Am I doing this right?? Use the second derivative test to locate all relative extrema and inflection points of f(x)= 2x^4-8x+3

OpenStudy (anonymous):

f'(x)= 8x^3-8 f"(x)=24x^2 24x^2=0 sqrt 24 = sqrt x x= 4.9

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

@phi what is the next step here to solve the problem?

OpenStudy (phi):

to find extremes, solve f'(x)=0 8x^3-8 = 0 x^3 = 1 x = 1 (and 2 complex roots) to find what type of extrema this is, use the 2nd derivative. f''(x)= 24x^2 at x=1 you get f''(1)= 24. It is positive, which means you found a minimum

OpenStudy (anonymous):

ok got it . thanks!

OpenStudy (anonymous):

@phi I'm sorry, one more question? how do i get the inflection points again?

OpenStudy (phi):

inflection points, max pts and min pts all have 0 slope i.e. 1st derivative is zero at those types of points. here there is only 1 point where f'(x) =0. It occurs at x=1. According to the 2nd derivative test, it is a minimum

OpenStudy (anonymous):

So The only inflection point is x=1

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

x=1 is not an inflection point. x=1 is a minimum see http://mathworld.wolfram.com/InflectionPoint.html